For each equation below, solve for X. 1. 1 +7c = 1 2x 2. 3.54y2 - V4x = 0 3. x2 (y-At) 4 4. 0.00753x + x = 0.023 5. 3.020x

please write legible.
Answer #1

22 or, ex 2x OV 2x tte T.: Pantai Assuming ming et l.fc a, (3.54-7292 1. + Fe = 1 E1-70 2(1-70) x = 21-70) 2. 3.54y2- Sex =4. 0.00 753% + 4 Z = 0.023 Z3X 0.00753x + L Z = 0.023 [ to Assuming 2 +9 ai 0.00 753x = 0.023 - 42 2 x = -(0.023 - 43). 0.007Hopefully you understand my steps . If you still have any doubt about any step please let me know in comments . I shall actively reply you. Thank you.